OpenStudy (anonymous):

Graph the following function: y=csc(1/2x) find the domain, range, period, even/odd function, and asymptotes

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

is this \[\csc(\frac{1}{2x})\] or \[csc(\frac{1}{2}x)\]?

OpenStudy (anonymous):

the 2nd one

OpenStudy (anonymous):

can you help?

OpenStudy (anonymous):

yeah i hope

OpenStudy (anonymous):

\[\csc(\frac{x}{2})=\frac{1}{\sin(\frac{x}{2})}\]

OpenStudy (anonymous):

so range is from \[-\infty\] to \[\infty\]

OpenStudy (anonymous):

sine is odd which means \[\sin(-x)=-\sin(x)\] and therefore this one is odd too

OpenStudy (anonymous):

of course you cannot divide by zero, therefore we have to make sure that \[\sin(\frac{x}{2})\neq0\]

OpenStudy (anonymous):

\[sin(x)=0\] means \[x=...-3\pi,-2\pi,-\pi,0, \pi, 2\pi, 3\pi...\]

OpenStudy (anonymous):

we have \[sin(\frac{x}{2})=0\] giving \[\frac{x}{2}=k\pi\] and therefore \[x=2k\pi\] so domain is all real numbers except \[x = 2k\pi, k\in Z\]

OpenStudy (anonymous):

ok i made a mistake earlier when i said range was all real numbers. sine is bounded above by 1 and below by -1 so this function is never between -1 and 1. in other words the domain is \[(-\infty,-1)\cup (1,\infty)\]

OpenStudy (anonymous):

how is it so far?

OpenStudy (anonymous):

except of course for my mistake!

OpenStudy (anonymous):

I am taking notes as we go, thank you so much

OpenStudy (anonymous):

well don't write the range is all real numbers. delete that from your notes!

OpenStudy (anonymous):

ok thank you so range is ?

OpenStudy (anonymous):

\[(-\infty,-1]\cup [1,\infty)\]

OpenStudy (anonymous):

oh excuse me, i wrote that as the domain

OpenStudy (anonymous):

it is -1 when \[sin(\frac{x}{2})=-1\]and 1 when \[\sin(\frac{x}{2})=1\]

OpenStudy (anonymous):

otherwise \[-1<\sin(\frac{x}{2})<1\] so \[\frac{1}{\sin(\frac{x}{2})}>1\] or \[\frac{1}{\sin(\frac{x}{2})}<-1\]

OpenStudy (anonymous):

what is left? asymptotes. we have found those. the function is undefined at \[x=2k\pi, k\in Z\] and so those are the asymptotes

OpenStudy (anonymous):

and finally the period! we know that the period of sine is \[2\pi\] so the period of \[\frac{1}{\sin(x)}\] is also \[2\pi\] and therefore the period of \[\frac{1}{sin(\frac{x}{2})}\] is \[4\pi\]

OpenStudy (anonymous):

and that takes care of all the questions yes?

OpenStudy (anonymous):

unless of course you would like to graph it and see what it looks like

OpenStudy (anonymous):

yes, thank you so much, i have one problem left if you have timeo look at it

OpenStudy (anonymous):

go ahead,

OpenStudy (anonymous):

y= tan(\[\pi\]x)+1

OpenStudy (anonymous):

that is all suppose to be in a line

OpenStudy (anonymous):

that is all suppose to be in a line

OpenStudy (anonymous):

\[tan(\pi x)+1\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

ok this one is much easier

OpenStudy (anonymous):

tangent is odd, but this is not odd because of the 1 outside

OpenStudy (anonymous):

in other words if \[f(x)=\tan(x)+1\] then \[f(-x)=\tan(-x)+1=-\tan(x)+1\neq -f(x)\]

OpenStudy (anonymous):

this one does have range all real numbers because tangent has range all real numbers.

OpenStudy (anonymous):

domain we find as before; \[tan(\pi x)+1=\frac{sin(\pi x)}{cos(\pi x)}+1\] and since you cannot divide by 0 we have to make sure that \[\cos(\pi x)\neq 0\]

OpenStudy (anonymous):

\[\cos(x)=0\] means \[x=...-\frac{5\pi}{2},-\frac{3\pi}{2},-\frac{\pi}{2},\frac{\pi}{2}, \frac{3\pi}{2}, \frac{5\pi}{2}, ...\]

OpenStudy (anonymous):

so \[cos(\pi x)=0 \] means \[x=-\frac{5}{2}, -\frac{3}{2}, -\frac{1}{2},\frac{1}{2},\frac{3}{2},\frac{5}{2}...\]

OpenStudy (anonymous):

more succinctly written as \[x\neq \frac{2k+1}{2}, k\in Z\]

OpenStudy (anonymous):

those are you asymptotes as well

OpenStudy (anonymous):

and finally the period of tangent is \[\pi\] so the period of \[\tan(\pi x)\] is \[\frac{\pi}{\pi}=1\]

OpenStudy (anonymous):

and that is that. except for the graph which is here http://www.wolframalpha.com/input/?i=tan%28pi+x%29%2B1

OpenStudy (anonymous):

is this even or neither

OpenStudy (anonymous):

no not even or odd.

OpenStudy (anonymous):

tangent is odd, but the +1 at the end throws it off

OpenStudy (anonymous):

and domain is

OpenStudy (anonymous):

all real numbers except \[x=\frac{2k+1}{2}, k\in Z\]

OpenStudy (anonymous):

in other words all numbers except 1/2,3/2,5/2,...

OpenStudy (anonymous):

odd integers divided by 2

OpenStudy (anonymous):

ok and is there a way to look at the graph

OpenStudy (anonymous):

i sent it i think

OpenStudy (anonymous):

oh im soryy mine did not show a graph

OpenStudy (anonymous):

those vertical lines are just asymptotes not part of the graph. did ou see the second one?

OpenStudy (anonymous):

yes thank you is ther an email adress that i can contact you at when i need additional help

OpenStudy (anonymous):

you can try me at satellite73.openstudy@gmail.com

OpenStudy (anonymous):

thank you alot i really appreciate it